# MIPHEI-vit: Multiplex Immunofluorescence Prediction from H&E Images using ViT Foundation Models

### v1

@@ Abstract
Histopathological analysis is a cornerstone of cancer diagnosis, with Hematoxylin and Eosin (H&E) staining routinely acquired for every patient to visualize cell morphology and tissue architecture. On the other hand, multiplex immunofluorescence (mIF) enables more precise cell type identification via proteomic markers, but has yet to achieve widespread clinical adoption due to cost and logistical constraints. To bridge this gap, we introduce MIPHEI (Multiplex Immunofluorescence Prediction from H&E), a U-Net-inspired architecture that integrates state-of-the-art ViT foundation models as encoders to predict mIF signals from H&E images. MIPHEI targets a comprehensive panel of markers spanning nuclear content, immune lineages (T cells, B cells, myeloid), epithelium, stroma, vasculature, and proliferation. We train our model using the publicly available ORION dataset of restained H&E and mIF images from colorectal cancer tissue, and validate it on two independent datasets. MIPHEI achieves accurate cell-type classification from H&E alone, with F1 scores of 0.88 for Pan-CK, 0.57 for CD3e, 0.56 for SMA, 0.36 for CD68, and 0.30 for CD20, substantially outperforming both a state-of-the-art baseline and a random classifier for most markers. Our results indicate that our model effectively captures the complex relationships between nuclear morphologies in their tissue context, as visible in H&E images and molecular markers defining specific cell types. MIPHEI offers a promising step toward enabling cell-type-aware analysis of large-scale H&E datasets, in view of uncovering relationships between spatial cellular organization and patient outcomes.
@@ 1 Introduction
Multiplex immunofluorescence (mIF) Andreou et al. (2022); Tan et al. (2020); Im et al. (2019) imaging is a powerful technique that improves the analysis of tissue sections by providing detailed information beyond what H&E staining can reveal. mIF achieves this by simultaneously visualizing and quantifying multiple protein markers within a single tissue section, utilizing fluorescently labeled antibodies that bind to specific proteins, allowing for the identification of cell types based on marker expression. This capability makes mIF useful across various domains, including cancer biology, immunology, and infectious disease, where understanding spatial cell organization is critical. By combining molecular and spatial information at single-cell resolution, mIF supports detailed characterization of tissue microenvironments and cellular interactions.
The evolution of immunolabeling techniques from traditional immunohistochemistry (IHC) Duraiyan et al. (2012) has led to more advanced mIF techniques like PhenoCycler Black et al. (2021), which enables the detection of up to 100 markers through multiple imaging cycles, each capturing 4 channels. This iterative process can however degrade tissue integrity. Among recent advancements, the Orion scanner introduced by Lin et al. Lin et al. (2023a) allows the simultaneous detection of up to 20 markers in a single cycle, preserving tissue quality while still providing rich multiplexed information. Orion also allows capturing high-quality, restained H&E images from the same tissue section.
While mIF imaging offers several advantages, it also presents important challenges. Preparing and processing mIF slides is time-consuming and labor-intensive, requiring expensive reagents and specialized equipment, which are not always available in all clinical settings, limiting its accessibility. Compared to other techniques for spatial biology, such as Imaging Mass Cytometry (IMC) Mann et al. (2001) and VisiumHD spatial transcriptomics Oliveira et al. (2024), mIF detects fewer molecular markers but offers much higher spatial resolution. Like these technologies, mIF is constrained by high costs, which prevents it from being adopted in clinical practice.
@@
In this study, we aim to predict the expression of key markers from H&E, covering nuclear content (Hoechst), vasculature (CD31), immune populations (CD45, CD68, CD4, FOXP3, CD8a, CD45RO, CD20, PD-L1, CD3e, CD163), epithelial and stromal structures (E-cadherin, Pan-CK, SMA), and proliferation (Ki67). To achieve this, we introduce MIPHEI: Multiplex Immunofluorescence Prediction from H&E-stained Whole Slide Images, a U-Net-inspired architecture integrating state-of-the-art foundation models as encoders. Unlike traditional cell type classification models that rely on manual labels or pseudo-labels, our method is trained directly on aligned mIF data, avoiding biases from predefined annotations. It is also highly modular and does therefore not rely on specific nucleus segmentation and cell classification methods.
@@
  • Rigorous Validation and Reproducibility: We demonstrate generalization of the method through validation with cell-level metrics for accurate cell type identification on both internal and two external datasets, moving beyond pixel-level evaluation. Our study sets a new benchmark framework for H&E to mIF translation.
@@
  • ViT Foundation Models in U-Net Architecture: one major methodological novelty relies in the integration of Vision Transformer (ViT) Dosovitskiy et al. (2020) foundation models as the backbone in a U-Net architecture for histopathological image translation, leveraging advanced representation learning.
Figure 1: Preprocessing pipeline: (a) H&E and mIF images are aligned using Valis Gatenbee et al. (2023). Tissue regions are then selected via Otsu thresholding on H&E, and a trained CNN filters misaligned tiles caused by the restaining and acquisition process. (b) Autofluorescence is subtracted from mIF images, followed by DAPI-based nuclei segmentation and nuclei dilation to approximate cell boundaries. (c) Pseudo-labels are generated by computing per-cell mean marker expression and applying GMM clustering to define marker positivity, determining labels (e.g., CD3e+).
@@ 2.1 Virtual Staining
Image-to-image translation involves transforming images from one domain to another, with Pix2Pix Isola et al. (2017) being one of the most well-known methods in this domain for paired images. Utilizing conditional adversarial networks, these techniques enable high-quality image synthesis, supporting tasks such as style transfer, super-resolution, and domain adaptation, and have driven significant advancements across various applications, as demonstrated by Wang et al. Wang et al. (2018).
In recent years, image-to-image translation techniques have become increasingly popular for life science applications, such as predicting super-resolved microscopy images Wang et al. (2019) or bright-field-like images from holographic images Wu et al. (2019). In histopathology, such virtual staining techniques have been used to predict immunohistochemistry (IHC) data Sun et al. (2023); DoanNgan et al. (2022), to do virtual multiplexing Wu et al. (2023); Bao et al. (2021), to predict mIF images from IHC images Ghahremani et al. (2022), or both H&E Rivenson et al. (2019); Bai et al. (2023); Zhang et al. (2020); Cao et al. (2023) and mIF Christiansen et al. (2018) from unlabeled autofluorescence images.
In the scope of this study, the most relevant work involves translating H&E images to mIF, which directly addresses the challenge of performing cell type calling from standard histological stains. A notable example is SHIFT Burlingame et al. (2020) based on conditional generative adversarial network (cGAN) to predict markers like DAPI, pan-cytokeratin, and $\alpha$-smooth muscle actin. Another relevant work is ImmunoAIzer Bian et al. (2021) using a semi-supervised adversarial approach to predict CD3, CD20, PanCK, and DAPI channels from H&E, trained on both aligned and unaligned data. More recently, the same author published HEMIT Bian et al. (2024), releasing a dataset and a method designed for translating H&E to mIHC images targeting DAPI, CD3, and pan-cytokeratin (Pan-CK) markers. The authors propose a hybrid ViT-CNN generator architecture combining CNNs with Swin Transformers Liu et al. (2021).
@@ 2.2 ViT on Dense Prediction Tasks
Figure 2: MIPHEI architecture: A U-Net-inspired model, based on VitMatte, is trained to predict mIF images from H&E, using the H-optimus-0 ViT foundation model, Tanh activation, and a custom weighted MSE loss to correct for the unbalanced distribution across markers.
ViTs Dosovitskiy et al. (2020) are now widely used in image classification and often outperform CNNs when pretrained with advanced contrastive self-supervised learning (SSL) methods such as MoCov3 Chen et al. (2021), iBOT Zhou et al. (2021), and DINOv2 Oquab et al. (2023). These SSL approaches have been adapted to histopathology, enabling recent foundation models to achieve strong and robust performance across diverse downstream tasks. CTransPath Wang et al. (2022) uses MoCov3 Chen et al. (2021) with a Swin Transfomer architecture Liu et al. (2021), and was trained on 15.6 million tiles from opensource datasets. UNIv2 Chen et al. (2024) leverages DINOv2 to train a ViT-H/14 model on 200 million tiles from 3.5k proprietary H&E and IHC slides. H-optimus-0 Saillard et al. (2024) is a ViT-G/14 variant with SSL pretraining based on iBOT Zhou et al. (2021) and DINOv2 Oquab et al. (2023), trained on hundreds of millions of tiles from 500,000+ H&E whole slide images. These models are among the state-of-the-art in computational pathology, with DINOv2-based models like UNIv2 and H-optimus-0 outperforming ImageNet-pretrained encoders and CTransPath on unseen datasets.
While plain ViTs are effective for image-level tasks, they often struggle with dense prediction tasks like image translation due to limited ability to capture fine local details, unlike CNNs which excel at leveraging local continuity and multiscale features. To overcome this, hybrid CNN-ViT architectures have been proposed. CellViT Hörst et al. (2024), for example, uses the UNETR architecture Hatamizadeh et al. (2022), which integrates a ViT encoder into a U-Net-like design, employing convolutional transpose blocks to create hierarchical features, offering a simple adaptation. ViTMatte Yao et al. (2024) combines a plain ViT with a convolutional module for pyramidal feature extraction and detail refinement, using ViT token features only as the bottleneck. More advanced models like ViT-Adapter Chen et al. (2022) and ViT-CoMer Xia et al. (2024) go further by enabling richer multi-scale interactions between ViT and CNN components, but at higher computational cost. While these hybrid methods have mainly been applied to segmentation, we hypothesized they could also benefit dense image translation tasks like mIF prediction, especially when using foundation models trained on large, diverse datasets.
@@ 3 Dataset
We present the datasets used in this study, with Figure 3 providing additional details on tissue distribution, domain shifts, and dataset composition. The preprocessed data for ORION and HEMIT datasets is available on Zenodo.
@@ 3.1 Orion CRC Dataset (Train/Val/Test)
The ORION colorectal cancer (CRC) dataset Lin et al. (2023b, a) was acquired with a novel system capturing 18-channel immunofluorescence (IF) images alongside H&E staining on the same tissue samples. This dataset includes 41 Whole Slides Images (WSIs) with both H&E and mIF data. The 15 markers (with additional DNA stain) of interest for this study are: Hoechst, CD31, CD45, CD68, CD4, FOXP3, CD8a, CD45RO, CD20, PD-L1, CD3e, CD163, E-cadherin, Pan-CK, SMA and Ki67. The PD-1 channel was not used due to poor signal quality. The H&E images were acquired with an Aperio GT450 microscope (Leica Biosystems) and registered to the IF images in the original study, at a resolution of 0.325 microns per pixel (mpp).
@@ 3.2 HEMIT Dataset (Test)
The HEMIT dataset Bian et al. (2024) consists of H&E and multiplex-immunohistochemistry (mIHC) images the same tissue sections, acquired using Mantra system scanner (PerkinElmer, Waltham, MA, USA) with DAPI, CD3, and Pan-CK markers. It includes 5,292 paired 1024×1024-pixel patches aligned at pixel level at 40× magnification.
@@ 3.3 IMMUcan CRC Dataset (Test)
IMMUcan (Integrated iMMUnoprofiling of large adaptive CANcer patient cohorts) Hong et al. (2020) is a European initiative launched in 2019 to advance Tumor Micro Environment (TME) profiling. We use only the CRC cases, comprising 35 registered H&E and mIF slides from two cohorts, which include both advanced and less severe stages. H&E slides were scanned with a Hamamatsu NanoZoomer 2.0-HT (0.5 mpp), and mIF images were acquired on a PerkinElmer Vectra Polaris. The mIF panel includes DAPI, CD3, CD8, CD4, FOXP3, and Pan-CK. Unlike HEMIT and ORION, the H&E and mIF slides in IMMUcan come from consecutive sections, requiring a dedicated evaluation protocol detailed later in Metric Overview. Although currently private, the dataset is expected to become publicly available.
Figure 3: Dataset Overview: (a) Distribution of tissue samples and tiles across datasets and splits. (b) 2D UMAP visualization of H-Optimized-0 embeddings, highlighting domain shifts across our datasets. (c) Normalized cell type distribution across datasets and splits.

### v2

@@ Abstract
Histopathological analysis is a cornerstone of cancer diagnosis, with Hematoxylin and Eosin (H&E) staining routinely acquired for every patient to visualize cell morphology and tissue architecture. On the other hand, multiplex immunofluorescence (mIF) enables more precise cell type identification via proteomic markers, but has yet to achieve widespread clinical adoption due to cost and logistical constraints. To bridge this gap, we introduce MIPHEI (Multiplex Immunofluorescence Prediction from H&E Images), a U-Net-inspired architecture that leverages a ViT pathology foundation model as encoder to predict mIF signals from H&E images using rich pretrained representations. MIPHEI targets a comprehensive panel of markers spanning nuclear content, immune lineages (T cells, B cells, myeloid), epithelium, stroma, vasculature, and proliferation. We train our model using the publicly available OrionCRC dataset of restained H&E and mIF images from colorectal cancer tissue, and validate it on five independent datasets: HEMIT, PathoCell, IMMUcan, Lizard and PanNuke. On OrionCRC test set, MIPHEI achieves accurate cell-type classification from H&E alone, with F1 scores of 0.93 for Pan-CK, 0.83 for $\alpha$-SMA, 0.68 for CD3e, 0.36 for CD20, and 0.28 for CD68, substantially outperforming both a state-of-the-art baseline and a random classifier for most markers. Our results indicate that, for some molecular markers, our model captures the complex relationships between nuclear morphologies in their tissue context, as visible in H&E images and molecular markers defining specific cell types. MIPHEI offers a promising step toward enabling cell-type-aware analysis of large-scale H&E datasets, in view of uncovering relationships between spatial cellular organization and patient outcomes.
Published in Computers in Biology and Medicine (2026), Open Access.
DOI: 10.1016/j.compbiomed.2026.111564
© 2026. Licensed under CC BY-NC-ND 4.0: creativecommons.org/licenses/by-nc-nd/4.0
Preprint: 15 May 2025 | Accepted: 19 February 2026
@@ 1 Introduction
Multiplex immunofluorescence (mIF) (Andreou et al., 2022; Tan et al., 2020; Im et al., 2019) imaging is a powerful technique that improves the analysis of tissue sections by providing detailed information beyond what H&E staining can reveal. mIF achieves this by simultaneously visualizing and quantifying multiple protein markers within a single tissue section, utilizing fluorescently labeled antibodies that bind to specific proteins, allowing for the identification of cell types based on marker expression. This capability makes mIF useful across various domains, including cancer biology, immunology, and infectious disease, where understanding spatial cell organization is critical. By combining molecular and spatial information at single-cell resolution, mIF supports detailed characterization of tissue microenvironments and cellular interactions.
The evolution of immunolabeling techniques from traditional immunohistochemistry (IHC) (Duraiyan et al., 2012) has led to more advanced mIF techniques like PhenoCycler (Black et al., 2021), which enables the detection of up to 100 markers through multiple imaging cycles, each capturing 4 channels. This iterative process can however degrade tissue integrity. Among recent advancements, the ORION scanner introduced by Lin et al. (2023b) allows the simultaneous detection of up to 20 markers in a single cycle, preserving tissue quality while still providing rich multiplexed information. ORION also allows capturing high-quality, restained H&E images from the same tissue section.
While mIF imaging offers several advantages, it also presents important challenges. Preparing and processing mIF slides is time-consuming and labor-intensive, requiring expensive reagents and specialized equipment, which are not always available in all clinical settings, limiting its accessibility. Compared to other techniques for spatial biology, such as Imaging Mass Cytometry (IMC) (Mann et al., 2001) and VisiumHD spatial transcriptomics (Oliveira et al., 2024), mIF detects fewer molecular markers but offers much higher spatial resolution. Like these technologies, mIF is constrained by high costs, which prevents it from being adopted in clinical practice.
@@
In this study, we aim to predict the expression of key markers from H&E, covering nuclear content (Hoechst), vasculature (CD31), immune populations (CD45, CD68, CD4, FOXP3, CD8a, CD45RO, CD20, PD-L1, CD3e, CD163), epithelial and stromal structures (E-cadherin, Pan-CK, $\alpha$-SMA), and proliferation (Ki67). To achieve this, we introduce MIPHEI: Multiplex Immunofluorescence Prediction from H&E-stained Whole Slide Images, a U-Net-inspired architecture integrating state-of-the-art foundation models as encoders. Unlike traditional cell type classification models that rely on manual labels or pseudo-labels, our method is trained directly on aligned mIF data, avoiding biases from predefined annotations. It is also highly modular and does therefore not rely on specific nucleus segmentation and cell classification methods. An overview of the complete MIPHEI workflow is shown in Figure 1.
@@
  • ViT Foundation Models in U-Net Architecture: one major methodological novelty is the integration of pathology Vision Transformer (ViT) (Dosovitskiy et al., 2020) foundation models as the encoder in a U-Net architecture for histopathological image translation. This design benefits from rich representations learned from millions of diverse tiles across various staining protocols and sample preparations.
@@
  • Rigorous Validation and Reproducibility: We demonstrate generalization of the method through validation with cell-level metrics for accurate cell type identification on both internal and five external datasets, moving beyond pixel-level evaluation. We further provide a unified benchmark enabling direct comparison of multiple existing H&E-to-mIF methods.
@@
  • Open Benchmark and Code: We release the full pipeline, pretrained models, and datasets to support reproducibility and future research in H&E to mIF prediction, available on GitHub [^1] GitHub: https://github.com/Sanofi-Public/MIPHEI-ViT and Zenodo (Balezo et al., 2025), with an interactive demo on Hugging Face [^2] Demo: https://huggingface.co/spaces/Estabousi/MIPHEI-vit-demo (see Supplementary Video 1 for a demonstration of the interactive interface).
Figure 1: Overall workflow of the study. (A) Datasets used for training and evaluation. (B) Data preprocessing, including alignment, tiling, mIF preprocessing, and cell-type extraction. (C) Our proposed MIPHEI architecture for H&E-to-mIF prediction. (D) Pixel- and cell-level evaluation on in-domain and out-of-domain datasets.
@@ 2.1 Virtual Staining
Image-to-image translation involves transforming images from one domain to another, with Pix2Pix (Isola et al., 2017) being one of the most well-known methods in this domain for paired images. Utilizing conditional adversarial networks, these techniques enable high-quality image synthesis, supporting tasks such as style transfer, super-resolution, and domain adaptation, and have driven significant advancements across various applications, as demonstrated by Wang et al. (2018). More recently, diffusion-based generative models have gained significant attention for high-quality image generation, including models such as Stable Diffusion (Rombach et al., 2022).
In recent years, image-to-image translation techniques have become increasingly popular for life science applications, such as predicting super-resolved microscopy images (Wang et al., 2019) or bright-field-like images from holographic images (Wu et al., 2019). In histopathology, such virtual staining techniques have been used to predict immunohistochemistry (IHC) data (Sun et al., 2023; Bai et al., 2022), to do virtual multiplexing (Wu et al., 2023; Bao et al., 2021; Ounissi et al., 2025), to predict mIF images from IHC images (Ghahremani et al., 2022), or both H&E (Rivenson et al., 2019; Bai et al., 2023; Zhang et al., 2020; Cao et al., 2023) and mIF (Christiansen et al., 2018; Zhou et al., 2024; Loo et al., 2025) from label-free imaging, such as autofluorescence or phase contrast microscopy.
In the scope of this study, the most relevant work involves translating H&E images to mIF, which directly addresses the challenge of performing cell type calling from standard histological stains. Several approaches address this problem under different supervision regimes. Histoplexer (Andani et al., 2025) predicts proteomic signals from H&E using consecutive IMC tissue sections. Other methods instead rely on same-section H&E–mIF pairs, which avoids imperfect registration and small variations in cellular content that can arise with consecutive sections, and therefore facilitates reliable cell-level supervision. A notable example is SHIFT (Burlingame et al., 2020) based on conditional generative adversarial network (cGAN) to predict markers like DAPI, pan-cytokeratin (Pan-CK), and $\alpha$-smooth muscle actin. HEMIT (Bian et al., 2024, 2021) further introduces a public dataset together with a hybrid Convolutional Neural Network (CNN)–Transformer generator combining convolutional layers with Swin Transformers (Liu et al., 2021) for H&E-to-mIHC translation, targeting DAPI, CD3, and Pan-CK markers. More recently, diffusion-based approaches have been explored, notably DiffusionFT (Oh et al., 2025), which builds on the Stable Diffusion (Rombach et al., 2022) pipeline to enable efficient per-marker mIF prediction. Finally, beyond pixel-level image translation, ROSIE (Wu et al., 2025) uses a ConvNeXt (Woo et al., 2023) model to predict per-channel mean mIF signals within a small central region of H&E images, yielding downsampled virtual staining prediction when applied in a sliding-window manner.
@@ 2.2 ViT on Dense Prediction Tasks
ViTs (Dosovitskiy et al., 2020) are now widely used in image classification and often outperform CNNs when pretrained with advanced contrastive self-supervised learning (SSL) methods such as MoCov3 (Chen et al., 2021), iBOT (Zhou et al., 2021), and DINOv2 (Oquab et al., 2023). These SSL approaches have been adapted to histopathology, enabling recent foundation models to achieve strong and robust performance across diverse downstream tasks. CTransPath (Wang et al., 2022) uses MoCov3 (Chen et al., 2021) with a Swin Transfomer architecture (Liu et al., 2021), and was trained on 15.6 million tiles from opensource datasets. UNIv2 (Chen et al., 2024) leverages DINOv2 to train a ViT-H/14 model on 200 million tiles from 3.5k proprietary H&E and IHC slides. H-optimus-0 (Saillard et al., 2024) is a ViT-G/14 variant also trained with DINOv2 on hundreds of millions of tiles from 500,000+ H&E whole slide images (WSIs). These models are among the state-of-the-art in computational pathology, with DINOv2-based models like UNIv2 and H-optimus-0 outperforming ImageNet-pretrained encoders and CTransPath on unseen datasets.
While plain ViTs are effective for image-level tasks, they often struggle with dense prediction tasks like image translation due to limited ability to capture fine local details, unlike CNNs which excel at leveraging local continuity and multiscale features. To overcome this, hybrid CNN-ViT architectures have been proposed. CellViT (Hörst et al., 2024), for example, uses the UNETR architecture (Hatamizadeh et al., 2022), which integrates a ViT encoder into a U-Net-like design, employing convolutional transpose blocks to create hierarchical features, offering a simple adaptation. ViTMatte (Yao et al., 2024) combines a plain ViT with a convolutional module for pyramidal feature extraction and detail refinement, using ViT token features only as the bottleneck. More advanced models like ViT-Adapter (Chen et al., 2022) and ViT-CoMer (Xia et al., 2024) go further by enabling richer multi-scale interactions between ViT and CNN components, but at higher computational cost. While these hybrid methods have mainly been applied to segmentation, we hypothesized they could also benefit dense image translation tasks like mIF prediction, especially when using foundation models trained on large, diverse datasets.
@@ 3 Dataset
Table 1: Overview of the datasets used in this study. MPP denotes the resolution in microns per pixel. HEMIT images were downsampled by a factor of 2 to match the 20x resolution used in our experiments.
Dataset Split #Tiles Size (px) Scanner MPP
Same-section H&E–mIF datasets
OrionCRC Train 295096 $\times333$ Aperio GT450 0.5
Val 12402 $\times333$ Aperio GT450 0.5
Test 10952 $\times333$ Aperio GT450 0.5
HEMIT Test 5292 $\times512$ Mantra $\approx$0.5
PathoCell Test 4807 $\times256$ Pannoramic P250 0.377
Consecutive-section H&E–mIF datasets
IMMUcanCRC Test 16933 $\times1024$ NanoZoomer 2.0-HT 0.5
H&E nuclei datasets
PanNuke Test 7901 $\times256$ Multi $\approx$0.25
Lizard Test 4777 $\times256$ Multi $\approx$0.5
Figure 2: Cell-type distribution across datasets. Proportions of cell types across datasets and splits. Datasets highlighted in red are used for training and validation. Cell-type labels vary between datasets, and rare cell types may appear at very low proportions.
We present the datasets used in this study, with Table 1 summarizing their composition and acquisition details, and Figure 2 illustrating cell type distributions across datasets. The preprocessed data for OrionCRC and HEMIT datasets is available on Zenodo.
@@ 3.1 OrionCRC Dataset (Train/Val/Test)
The ORION colorectal cancer (OrionCRC) dataset (Lin et al., 2023a, b) was acquired with a novel system capturing 18-channel immunofluorescence (IF) images alongside H&E staining on the same tissue sections. This dataset includes 41 WSIs with both H&E and mIF data. The 15 markers (with additional DNA stain) of interest for this study are: Hoechst, CD31, CD45, CD68, CD4, FOXP3, CD8a, CD45RO, CD20, PD-L1, CD3e, CD163, E-cadherin, Pan-CK, $\alpha$-SMA and Ki67. The PD-1 channel was not used due to poor signal quality.
@@ 3.2 HEMIT Dataset (Test)
The HEMIT dataset (Bian et al., 2024) consists of H&E and multiplex-immunohistochemistry (mIHC) images the same tissue sections, with DAPI, CD3, and Pan-CK markers.
3.3 PathoCell Dataset (Test)
The PathoCell dataset was introduced in Schürch et al. (2020) and comprises 106 tissue microarrays (TMAs) cores obtained from around 70 colorectal cancer (CRC) specimens originating from 35 patients. Tissues were imaged using a 56-marker PhenoCycler (Black et al., 2021) mIF panel together with same section H&E. The dataset was later adopted in Lüscher et al. (2025) to evaluate foundation models for cell phenotyping. In this setting, cell instances and cell phenotypes were derived from the mIF measurements (cf. Section 3.6.4), resulting in 15 categories. For this dataset, we applied only minimal preprocessing (cf Section 3.6.4), providing a clearly independent dataset for evaluation.
@@ 3.4 IMMUcan CRC Dataset (Test)
IMMUcan (Integrated iMMUnoprofiling of large adaptive CANcer patient cohorts) (Hong et al., 2020) is a European initiative launched in 2019 to advance Tumor Micro Environment (TME) profiling. We use only the CRC cases, comprising 35 registered H&E and mIF WSIs from two cohorts, which include both advanced and less severe stages. mIF images were acquired on a PerkinElmer Vectra Polaris and the mIF panel includes DAPI, CD3, CD8, CD4, FOXP3, and Pan-CK. Unlike HEMIT, PathoCell and OrionCRC, the H&E and mIF slides in IMMUcan come from consecutive sections, requiring a dedicated evaluation protocol detailed later in Section 5.2.2. Although currently private, the dataset is expected to become publicly available.
3.5 Expert-annotated H&E nuclei datasets (Test)
To further evaluate MIPHEI at the cell level, we use two expert-annotated H&E nuclei datasets: Lizard (Graham et al., 2021) and PanNuke (Gamper et al., 2020). These datasets provide ground-truth cell-type annotations directly from H&E images and are used exclusively for out-of-domain evaluation. Lizard is a multi-center CRC cohort containing 291 fields of view. PanNuke spans 18 organs and provides expert annotations of nuclei into five cell categories. It is the only dataset in our evaluation that includes non-CRC tissue. For both datasets, cell-type labels are derived from expert annotations on H&E images and therefore correspond to coarser categories than mIF-derived cell phenotypes.